Suppose that f:[0,2]→ R is continuous on [0,2] and differentiable on [0,2] and


that f(0) =0 , f(1) =1, f(2) =1.


(i) Show that there exists c↓1∈ (0,1)such that f'(c↓1) =1


(ii) Show that there exists c↓2 ∈ (1,2)such that f'(c↓2) =0.


(iii) Show that there exists c ∈ (0,2)such that f'(c) =1/3



1
Expert's answer
2022-03-21T15:09:02-0400

"\\left( i \\right) Lagrange\u0091s\\,\\,theorem\\,\\,on\\,\\,\\left[ 0,1 \\right] :\\\\f\\left( 1 \\right) -f\\left( 0 \\right) =f'\\left( \\xi \\right) \\left( 1-0 \\right) \\Rightarrow f'\\left( \\xi \\right) =1-0=1\\\\\\left( ii \\right) Lagrange\u0092s\\,\\,theorem\\,\\,on\\,\\,\\left[ 1,2 \\right] :\\\\f\\left( 2 \\right) -f\\left( 1 \\right) =f'\\left( \\xi \\right) \\left( 2-1 \\right) \\Rightarrow f'\\left( \\xi \\right) =1-1=0\\\\\\left( iii \\right) f'\\in C\\left[ 0,2 \\right] ,f'\\left( \\xi _1 \\right) =0,f'\\left( \\xi _2 \\right) =1\\Rightarrow \\\\\\Rightarrow by\\,\\,IMVT\\,\\,there\\,\\,exists\\,\\,\\xi \\,\\,between\\,\\,\\xi _1,\\xi _2: f'\\left( \\xi \\right) =\\frac{1}{3}"